A newly issued 20-year maturity, zero-coupon bond is issued with a yield to maturity of 8% and face value of $1,000. If the applicable tax rate is 21% and you own 10 of these bonds, what will be your tax liabilities for the next two years?
Answer #1

Value of Bond today = Maturity Value / (1+r)^n

= 1000 / (1+0.08)^20

= 214.55

Value of Bond in 2 yrs = Maturity Value / (1+r)^n

= 1000 / (1+0.08)^19

= 231.71

Value of Bond today = Maturity Value / (1+r)^n

= 1000 / (1+0.08)^18

= 250.25

Tax in year 1 = (Value in year 1 - Value today) * no of Bonds * Tax

= (231.71 - 214.55) * 10 * 21%

= 36.04

Tax in year 2 = (Value in year 2-Value in year 2) * no of Bonds * Tax

= (250.25 - 231.71) * 10 * 21%

= 38.93

Since no interest is paid on Zero Coupon bond. but the Capital Gain attracts tax. So Tax is paid on the Capital Gain Appretiation.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.